In exercises 38-41 write each measurement in the specified unit.
38) 4 m; cm
40) 108 ft; yd

Answer by stanbon(75887)

write each measurement in the specified unit.
38) 4 m; cm
= 4m(100cm/1m) = 4*100 cm = 400cm
======================================
40) 108 ft; yd
108 ft(1 yd/3 ft) = (108/3)yd = 36 yd
==========================================
Cheers,
Stan H.